# Mastering Collaborative Multi-modal Data Selection: A Focus on Informativeness, Uniqueness, and Representativeness

### v1

@@ 3.4 Representative Value in Inter-cluster Space
Although the unique value effectively enhances the uniqueness of selected data within clusters, it overlooks their representativeness across the overall dataset, potentially allowing outlier noisy data to affect selection. Empirically, Clusters evaluated solely for uniqueness may overlook outlier noisy samples, which exhibit weaker associations with other clusters and limit the ability of selected data to represent the overall distribution patterns. Therefore, we introduce an inter-cluster association coefficient to measure relationships across clusters, ensuring that selected samples capture representative features from the overall dataset, thereby avoiding the selection of noisy data:
$\tau_{i}^{c}=\frac{1}{K-1}\sum_{k\neq c}^{K}\exp({sim}(\mathbf{\overline{p_{k}}},\mathbf{\overline{p_{c}}}))$ (4)
where $\mathbf{\overline{p_{c}}}$ is the average latent feature of all instructions in the target cluster $\mathbf{C}$ contains instruction $s_{i}$ and $\{\mathbf{\overline{p_{k}}}\}_{k\neq c}^{K}$ is the average feature of other clusters in the specific task. We use the feature of the last token to represent each instruction and $sim(\cdot,\cdot)$ denotes the cosine similarity. Based on this coefficient, we then assign the weighted representative value to the instruction $s_{i}$ as follows:
$V_{i}^{Rep}=\tau^{c}_{i}\cdot V_{i}^{Inf}$ (5)
In this way, the representative value uses the association coefficient to ensure that selected samples align with the overall distribution. When the value is high, it indicates that the selected samples can effectively represent other samples, reducing the impact of noisy data and enhancing the overall representativeness in conjunction with uniqueness.
@@ 3.5 Adaptively Collaborative Data Selection
Although we obtain multi-scale values of multi-modal instructions from three complementary perspectives, combining them to select ideal samples is challenging due to inconsistencies in sample structure within the candidate dataset. Specifically, multi-turn instructions should prioritize informative value due to their weak interrelationships, whereas single-turn instructions should emphasize unique and representative value due to their limited internal information. Inspired by this, we introduce an influence factor based on the number of response rounds of each multi-modal instruction for adaptively collaborative data selection to enable adaptive, collaborative data selection, enhancing the synergy among these three values as follows:
$V_{i}=\frac{N_{i}}{N_{i}+2}\cdot{V}_{i}^{Inf}+\frac{1}{N_{i}+2}\cdot({V}_{i}^{Uni}+{V}_{i}^{Rep})$ (6)
where $N_{i}$ denotes the conversation round of each multi-modal instruction. With this synergistic value for data selection, we can identify informative and unique instructions while also being adequately representative (c.f., Fig. 4).
In addition, we observe that standardizing the data selection proportion across all tasks limits selection diversity due to differences in task difficulty. To address this, we propose adaptively determining the data selection proportion for each task based on the average largest singular value ratio, which correlates with training convergence and reflects task difficulty. The data selection proportion $k_{p}$ for each task $S_{p}$ is computed as follows:
$k_{p}=\frac{x_{p}^{2}\cdot|S_{p}|}{\sum_{q}x_{q}^{2}\cdot|S_{q}|}\cdot k,\quad x_{p}=\overline{\frac{\sigma_{0}}{\sum_{j=1}^{L_{i}}\sigma_{j}}}$ (7)
where $x_{q}$ is the average of the largest singular value ratios for all samples in the task $S_{q}$, $|S_{q}|$ is the corresponding number of its samples. According to the above formula, the data selection rate of each task is adjusted from $k$ to $k_{p}$ to achieve task-adaptive proportions. Through collaborative value assessment with task-adaptive proportions, DataTailor promotes more diversity during MLLM data selection. More details and analyses are shown in Appendix B.2.
@@
MME-P $\uparrow$ MME-C $\uparrow$ SEED-Bench-I $\uparrow$ POPE $\uparrow$ VizWiz $\uparrow$ ScienceQA $\uparrow$ GQA $\uparrow$ VQA-v2$\uparrow$ TextVQA $\uparrow$ NoCaps (val) $\uparrow$
MiniGPT4-Instruction Model: MiniGPT4-7B
MiniGPT4-7B 3.4k 717.37 259.55 23.8 68.3 36.0 36.3 32.2 32.1 21.4 111.5
   Random 0.2k 698.43 227.66 25.1 69.7 18.3 34.0 19.2 33.2 17.2 105.1
   Length 0.2k 683.39 209.55 26.7 69.8 29.9 35.6 32.5 33.7 17.4 106.5
   E2LN [42] 0.2k 668.55 207.64 26.5 72.0 41.9 36.1 32.9 36.3 23.7 108.3
   IFD [30] 0.2k 678.61 213.75 29.1 47.4 42.7 38.1 28.3 36.0 23.4 106.6
   InsTag [37] 0.2k 715.64 237.86 26.8 70.4 40.0 38.1 30.1 34.5 22.2 105.9
   LESS [54] 0.2k 698.47 191.36 22.4 71.8 38.4 35.4 26.0 34.4 16.6 109.7
   InstructionGPT-4 [51] 0.2k 716.94 229.64 17.4 71.6 29.9 35.1 26.8 34.8 22.1 106.8
   SELF-FILTER [53] 0.5k 438.73 128.57 21.7 71.4 41.3 35.7 30.4 35.0 22.0 105.6
   TIVE [36] 0.2k 707.02 200.86 23.6 72.3 31.4 33.8 26.4 35.1 17.5 108.9
 DataTailor (Ours) 0.2k 720.63 263.93 27.3 69.8 40.8 37.7 30.7 34.7 21.0 106.9
LLaVA-1.5-mix-665k Model: LLaVA-7B
LLaVA-v1.5-7B (LoRA) 665k 1476.90 267.90 67.4 86.4 47.8 70.0 63.0 79.1 58.2 106.5
   Random 50k 1387.45 287.50 59.7 85.7 42.3 70.0 55.0 73.7 53.1 107.7
   Length 50k 1356.96 265.71 47.0 82.6 49.2 60.9 55.5 70.7 45.2 88.2
   E2LN [42] 50k 1077.31 252.50 59.3 80.8 44.4 71.0 41.7 61.0 41.7 86.9
   GradN [42] 50k 1275.44 303.57 58.3 75.7 37.8 70.9 44.9 64.0 46.0 101.9
   IFD [30] 50k 1113.44 301.79 55.1 76.7 48.7 48.2 41.9 64.2 43.6 106.8
   InsTag [37] 50k 1317.14 345.00 57.4 82.1 47.4 69.3 52.5 63.2 53.3 108.3
   LESS [54] 50k 1344.80 281.80 61.2 79.4 44.4 71.0 53.4 71.8 52.0 106.2
   SELF-FILTER [53] 25k 955.65 262.50 47.5 76.0 40.8 59.4 3.6 2.1 5.6 82.3
   TIVE [36] 50k 1334.80 248.57 62.2 85.9 45.1 71.4 56.2 73.8 51.1 96.0
 DataTailor (Ours) 50k 1461.23 362.50 61.7 82.1 46.3 70.9 57.7 75.0 53.1 107.2
 DataTailor w/ Increased Ratio (Ours) 100k 1476.15 319.29 63.6 85.3 49.5 71.0 60.5 76.7 55.7 108.7
@@ 4.1 Experimental Setup
Benchmarks. We assess our methods using a mix of general downstream tasks and MLLM-specific benchmarks, covering a wide range of capabilities. For general VQA tasks, VQA-v2 [5] and GQA [20] access the model’s visual perception abilities with open-ended questions while TextVQA [45] focuses on text-rich visual question answering. For general captioning tasks, we transfer MLLMs to NoCaps [2] validation set for zero-shot evaluation. Aligned with state-of-the-art (SOTA) MLLM methods, we include other benchmarks for comprehensiveness: MME [13] is used to evaluate MLLM’s reasoning ability from the two dimensions of perception and cognition; SEED-Bench [24] involves more comprehensive multi-modal tasks across 12 perspectives with the assistance of GPT-4, POPE [31] mainly evaluates the MLLM’s hallucination problems, VizWiz [16] and ScienceQA [44] contain unseen visual queries and multiple-choice questions to evaluate the ability of MLLMs to achieve zero-shot generalization from informative samples. We also present the corresponding tailored amount of valid data to demonstrate reduced training time.
Baselines. We use the following baselines: 1) Traditional data selection: it includes traditional random selection; length-based selection; GradN [42] and E2LN [42] use the L2-norm of the gradient and the error vector for selection, respectively. 2) LLM data selection: it directly transfers the method for selecting instruction data from LLMs to MLLMs, including heuristic score methods [7], human-reward methods [37], and gradient-based method [54]. 3) MLLM-specialized selection: it involves methods specifically designed for data selection in MLLM, including InstructionGPT-4 [51], SELF-FILTER [53], TIVE [36].
@@ 4.2 Main Results on Multi-modal Data Selection
Multi-modal instruction data suffers from serious redundancy, resulting in overall poor data quality. We can observe that in most cases, even randomly selecting a small amount of instruction data does not result in a performance drop proportional to the reduction in data size, particularly in general VQA benchmarks. Moreover, in some cases, simply selecting part of the data outperforms utilizing the entire dataset (33.3 v.s. 32.1 of VQA-v2 on MiniGPT-4), suggesting that excessive low-quality data hinder several MLLM’s capabilities on the contrary. Qualitatively, as shown in Fig. 1, most methods achieve 80% performance with less than 20% of the data, indicating that the additional data from the original MLLM does not significantly improve performance, but rather increases training time. This confirms our analysis of the data redundancy in multi-modal datasets and the necessity of data selection for MLLMs.
@@
For LLM data selection approaches ( i.e., IFD [30], InsTag [37], and LESS [54]), the performances across several benchmarks overall remain unsatisfactory. We notice that LESS significantly enhances the visual perception capabilities of MLLM, but it still struggles with complex comprehension questions (1344.8 of MME-P but only 281.8 of MME-C). A possible reason is that LESS prioritizes the most influential data without considering its uniqueness, reducing the selection of samples that contribute uniquely to reasoning capabilities. InsTag carefully curates unique data based on human rewards but is time-consuming and unscalable. Besides, due to the gap between pre-defined rules and sample values for MLLM fine-tuning, IFD demonstrates the poorest performance when directly transferring to MLLM data selection. Moreover, all LLM data selection methods demonstrate severe shortcomings in representativeness, which leads to a decline in the performance of general VQA tasks (average 49.6 in TextVQA and 49.3 in GQA). In contrast, We observe that DataTailor obtains overall improvement on various tasks. For an intuitive illustration of DataTailor paying attention to the informativeness, uniqueness, and representativeness of samples in data selection, we visually compare the corresponding values of these principles of DataTailor and those LLM data selection methods, as shown in Figure 1(b). Similarly, those LLM data selection methods exhibit deficiencies in corresponding dimensions, whereas DataTailor consistently achieves promising results from three perspectives. This result demonstrates DataTailor’s capability to effectively promote the collaboration of informative, unique, and representative values for multi-modal data selection, rather than roughly selecting samples based on individual values.
@@
Our DataTailor can be flexibly equipped to different MLLM for diverse multi-modal data selection. We incorporate our DataTailor into the two most popular multi-modal instruction datasets for evaluation, which contain both small-scale human-annotated instructions and large-scale data of various types. Despite the data diversity, our DataTailor can achieve consistently competitive performance across all benchmarks under limited data settings (e.g., 984.6 v.s. 976.9 on MME for MiniGPT4 with 5% data and 1823.7 v.s. 1744.8 on MME for LLaVA-v1.5 with 7.5% data). Notably, despite comprising only 7.5% data, DataTailor consistently outperforms full fine-tuning on the challenging zero-shot MLLM tasks (+3.9% in MiniGPT4 of SciQA and +1.3% in LLaVA-mix-665k). These results indicate that our proposed method effectively addresses data redundancy and consistently selects high-value samples across different categories of instruction data.
@@
Compared with other data selection methods, our DataTailor outperforms all of them in both MLLM benchmarks and downstream tasks. Specifically, DataTailor exceeds SOTA of MLLM-specialized selection methods [36, 51, 53] for all benchmarks with consistent improvements (average 99.7% of full performance v.s. 93.2% in TIVE [36]), especially +240.36 in MME compared with TIVE. Furthermore, we observe that SELF-FILTER [53] yields poor performance on downstream VQA tasks due to its reliance on pre-defined scoring networks, which limits the generalization of the samples that are selected by SELF-FILTER. Notably, when increases its ratio of data selection to 15%, DataTailor surpasses LLaVA-1.5’s full tuning (102.9% for MLLM performance and 100.8% for total performance). This indicates that data quality is significantly more crucial than large quantities of low-quality data for enhancing MLLMs. It truly exemplifies the characteristic of “Less is More” of DataTailor.
@@
mPLUG-Owl-7B [60]
    Full Data (100%) 1243.4 34.3 67.4 41.1 100.0%
    Random (5%) 1183.2 33.9 70.1 40.6 99.2%
    TIVE [36] (5%) 1177.1 34.0 70.3 41.3 99.6%
    DataTailor (5%) 1260.0 34.5 70.3 41.9 102.1%
Bunny-3B [17]
    Full Data (100%) 1778.1 70.7 86.8 70.9 100.0%
    Random (5%) 1578.8 64.4 83.4 70.1 93.7%
    TIVE [36] (5%) 1542.4 61.7 84.7 68.4 92.0%
    DataTailor (5%) 1582.8 65.3 81.4 75.8 95.6%

### v2

@@ 3.4 Representative Value in Inter-cluster Space
Empirically, evaluating samples solely with their informativeness and uniqueness within clusters may introduce outlier noisy or mislabeled data. These samples, despite higher uniqueness, have weaker associations with other clusters, limiting their ability to represent the overall dataset. Therefore, we introduce an inter-cluster representative coefficient to measure relationships across clusters, ensuring that selected samples capture representative features from the overall dataset, thereby avoiding the selection of noisy data:
$\tau_{i}^{c}=\frac{1}{K-1}\sum_{k\neq c}^{K}\exp({\operatorname{sim}}(\mathbf{\overline{p_{k}}},\mathbf{\overline{p_{c}}}))$ (5)
where $\mathbf{\overline{p_{c}}}$ is the average sample feature in the target cluster $\mathbf{C}$ that contains sample $s_{i}$, $\{\mathbf{\overline{p_{k}}}\}_{k\neq c}^{K}$ is the average sample feature of other clusters and $K$ is the number of clusters in each task. We use the feature of the last token to represent the sample feature as it aggregates all visual and textual features by cross-attention. Here, $\operatorname{sim}(\cdot,\cdot)$ is the cosine similarity and $\exp(\cdot)$ is used to amplify the effect of clusters. Based on this coefficient, we then assign the weighted representative value to the instruction $s_{i}$ as follows:
$V_{i}^{Rep}=\tau^{c}_{i}\cdot\frac{V_{i}^{Inf}}{\sum_{k\in\mathbf{C}}V_{k}^{Inf}}$ (6)
In this way, the representative value uses the association coefficient to ensure that selected samples align with the overall distribution. When the value is high, it indicates that the selected samples can effectively represent other samples, reducing the impact of noisy data and enhancing the overall representativeness in conjunction with uniqueness. More implementation details are shown in Appendix C.1.
@@ 3.5 Adaptively Collaborative Data Selection
Although we obtain multi-scale values from three complementary perspectives, combining them to select ideal samples is challenging since these multi-modal samples exhibit varying instruction rounds. Multi-IF [23] points out that multi-turn instructions prioritize informative value due to their weak interrelationships while single-turn ones emphasize unique and representative value due to their limited information. Inspired by this, we introduce an influence factor based on the number of response rounds of each multi-modal instruction for adaptively collaborative data selection to enable adaptive, collaborative data selection, enhancing the synergy among these three values as follows:
$V_{i}=\frac{r_{i}}{r_{i}+2}\cdot{V}_{i}^{Inf}+\frac{1}{r_{i}+2}\cdot({V}_{i}^{Uni}+{V}_{i}^{Rep})$ (7)
where $r_{i}$ denotes the conversation round of each multi-modal instruction. We employ a 1:1 ratio to simplify the balance between uniqueness and representativeness values, as their trade-off remains stable. With this synergistic value for data selection, we can identify informative and unique instructions while adequately representative (c.f., Fig. 4). We also show detailed demonstration of DataTailor addressing each of three principles in Sec. 4.4 and Appendix D.1.
In addition, we observe that standardizing the data selection proportion across all tasks limits selection diversity due to differences in task difficulty. Since spectral analysis shows that samples with the higher largest singular value exhibit less directional diversity for generalization, they require more selected data to improve the model’s training robustness. Thus, we propose adaptive data selection proportion $k_{p}$ for each task $S_{p}$ based on largest singular values, which is computed as follows:
$k_{p}=\frac{x_{p}^{2}\cdot|S_{p}|}{\sum_{q}x_{q}^{2}\cdot|S_{q}|}\cdot k,\quad x_{p}=avg({\frac{\sigma_{\text{max}}}{\sum_{j=1}^{L_{i}}\sigma_{j}}})$ (8)
where $x_{q}$ is the average ratio of the largest singular value $\sigma_{\text{max}}$ to the sum of all singular values for all samples in the task $S_{q}$, $|S_{q}|$ is the number of samples in the task, and $k$ is the whole data selection proportion. Through collaborative value assessment with task-adaptive proportions, DataTailor promotes more diversity during MLLM data selection. More details and analyses are shown in Appendix B.2.
@@
[…]
MiniGPT4-Instruction MiniGPT4-7B
MiniGPT4-7B 3.4k 717.4 259.6 23.8 68.3 19.0 25.1 23.4 36.0 36.3 32.2 32.1 21.4 111.5 100.0%
   Random 0.2k 698.4 227.7 25.1 69.7 17.2 27.9 19.1 18.3 34.0 19.2 33.2 17.2 105.1 89.1%
   Length 0.2k 683.4 209.6 26.7 69.8 18.3 26.7 17.4 29.9 35.6 32.5 33.7 17.4 106.5 94.7%
   EL2N [50] 0.2k 668.6 207.6 26.5 72.0 17.4 27.3 23.1 41.9 36.1 32.9 36.3 23.7 108.3 102.2%
   IFD [36] 0.2k 678.6 213.8 29.1 47.4 18.9 28.4 20.6 42.7 38.1 28.3 36.0 23.4 106.6 99.8%
   InsTag [44] 0.2k 715.6 237.9 26.8 70.4 15.8 28.1 21.7 40.0 38.1 30.1 34.5 22.2 105.9 100.8%
   LESS [62] 0.2k 698.5 191.4 22.4 71.8 19.8 25.8 22.2 38.4 35.4 26.0 34.4 16.6 109.7 95.4%
   InstructionGPT-4 [58] 0.2k 716.9 229.6 17.4 71.6 21.3 25.3 14.9 29.9 35.1 26.8 34.8 22.1 106.8 93.3%
   SELF-FILTER [60] 0.5k 438.7 128.6 21.7 71.4 19.4 24.3 20.4 41.3 35.7 30.4 35.0 22.0 105.6 92.8%
   TIVE [43] 0.2k 707.0 200.9 23.6 72.3 17.5 25.8 18.2 31.4 33.8 26.4 35.1 17.5 108.9 92.7%
 DataTailor (Ours) 0.2k 720.6 263.9 27.3 69.8 21.4 28.4 23.6 40.8 37.7 30.7 34.7 21.0 106.9 104.6%
LLaVA-1.5-mix-665k LLaVA-7B
LLaVA-v1.5-7B (LoRA) 665k 1476.9 267.9 67.4 86.4 30.9 67.9 32.8 47.8 70.0 63.0 79.1 58.2 106.5 100.0%
   Random 50k 1387.5 287.5 59.7 85.7 29.5 64.5 32.2 42.3 70.0 55.0 73.7 53.1 107.7 95.3%
   Length 50k 1357.0 265.7 47.0 82.6 29.7 67.6 33.9 49.2 60.9 55.5 70.7 45.2 88.2 91.0%
   EL2N [50] 50k 1077.3 252.5 59.3 80.8 21.1 40.1 33.6 44.4 71.0 41.7 61.0 41.7 86.9 82.3%
   GradN [50] 50k 1275.4 303.6 58.3 75.7 24.8 68.2 32.4 37.8 70.9 44.9 64.0 46.0 101.9 89.3%
   IFD [36] 50k 1113.4 301.8 55.1 76.7 27.6 63.1 33.0 48.7 48.2 41.9 64.2 43.6 106.8 87.3%
   InsTag [44] 50k 1317.1 345.0 57.4 82.1 29.6 68.1 34.0 47.4 69.3 52.5 63.2 53.3 108.3 96.4%
   LESS [62] 50k 1344.8 281.8 61.2 79.4 28.3 65.5 33.0 44.4 71.0 53.4 71.8 52.0 106.2 94.3%
   SELF-FILTER [60] 25k 955.7 262.5 47.5 76.0 26.6 60.3 30.6 40.8 59.4 3.6 2.1 5.6 82.3 65.8%
   TIVE [43] 50k 1334.8 248.6 62.2 85.9 30.2 67.9 33.1 45.1 71.4 56.2 73.8 51.1 96.0 94.6%
   COINCIDE [30] 133k 1495.6 - - 86.1 - 67.3 - 46.8 69.2 59.8 76.5 55.6 - 98.0%
   ICONS [61] 133k 1485.7 - - 87.5 29.7 66.1 - 50.1 70.8 60.7 76.3 55.6 - 98.8%
 DataTailor (Ours) 50k 1461.2 362.5 61.7 82.1 30.4 69.3 33.9 46.3 70.9 57.7 75.0 53.1 107.2 100.1%
 DataTailor (Ours) 100k 1476.2 319.3 63.6 85.3 31.8 71.1 33.2 49.5 71.0 60.5 76.7 55.7 108.7 101.3%
@@ 4.1 Experimental Setup
Benchmarks. We assess our methods using a mix of MLLM-specific benchmarks and more general downstream tasks. Aligned with SOTA MLLM methods, we include MLLM benchmarks for comprehensiveness: MME [17] is used to evaluate MLLM’s ability of perception and cognition; SEED-Bench [32] involves multi-modal tasks across 12 perspectives with the assistance of GPT-4; POPE [37] mainly evaluates the MLLM’s hallucination problems; MM-Vet [71] and LLaVA-Wild [42] assess the model’s open-ended conversational capabilities; The MMMU [72] consists of more challenging scientific problems to assess reasoning ability. For general VQA tasks, VizWiz [21] and ScienceQA [53] contain unseen visual queries and multiple-choice questions to evaluate the zero-shot generalization of MLLMs. VQA-v2 [5] and GQA [26] access the model’s visual perception abilities with open-ended questions while TextVQA [54] focuses on text-rich questions. For captioning, we transfer MLLMs to the NoCaps [2] validation set. We also show the amount of valid data selected to demonstrate the effectiveness of data selection. Note that Rel. in all tables represents the relative boost on all benchmarks
Baselines. We use the following baselines: 1) Traditional data selection: it includes traditional random selection; length-based selection; GradN [50] and EL2N [50] use the L2-norm of the gradient and the error vector for selection, respectively. 2) LLM data selection: it directly transfers the selection methods from LLMs to MLLMs, including data-specific methods [9], human-reward methods [44], and gradient-based method [62]. 3) MLLM-specialized selection: it involves methods specifically designed for data selection in MLLM, including InstructionGPT-4 [58], SELF-FILTER [60], TIVE [43], COINCIDE [30], ICONS [61].
@@ 4.2 Main Results on Multi-modal Data Selection
Multi-modal instruction data suffers from serious redundancy, resulting in overall poor data quality. We can observe that in most benchmarks, even randomly selecting a small amount of instruction data does not result in a performance drop. Moreover, in some cases, simply selecting part of the data outperforms utilizing the entire dataset (33.2 v.s. 32.1 of VQA-v2 on MiniGPT-4), suggesting that excessive low-quality data hinder several MLLMs’ capabilities on the contrary. Qualitatively, as shown in Fig. 1, most methods achieve 80% performance with less than 20% of the data. This confirms our analysis of the data redundancy in multi-modal datasets and the necessity of data selection.
@@
For LLM data selection approaches ( i.e., IFD [36], InsTag [44], and LESS [62]), the performances across several benchmarks overall remain unsatisfactory. Although these baselines explicitly distinguish tasks and features, they still underperform DataTailor by 12.8%, 3.7%, and 5.8% on overall relative performance due to their rough information modeling and disregard for sample relationships. Moreover, all LLM data selection methods demonstrate severe shortcomings in representativeness, which leads to a decline in the performance of general VQA tasks (average 49.6 in TextVQA and 49.3 in GQA). In contrast, we observe that DataTailor obtains overall improvement on various tasks. For an intuitive illustration, we visually compare the metrics of three principles of DataTailor and those LLM data selection methods, as shown in Figure 1(b). Similarly, those LLM data selection methods exhibit deficiencies in at least one dimension, whereas DataTailor consistently achieves promising results. This result demonstrates DataTailor’s capability to effectively promote these principles for multi-modal data selection, rather than roughly selecting samples based on individual values.
@@
Our DataTailor can be flexibly applied to different MLLMs. We incorporate our DataTailor into the two most popular MLLMs for evaluation. Despite the diversity in data and model structures, our DataTailor consistently improves relative performance of data selection across all benchmarks compared to random selection (e.g., +15.5% on MiniGPT4-7B and +4.8% on LLaVA-v1.5-7B). Notably, despite comprising only 7.5% data, DataTailor consistently outperforms full fine-tuning on the challenging MMMU benchmark (23.6 v.s. 23.4 on MiniGPT4-7B and 33.9 v.s. 32.8 on LLaVA-7B). These results indicate that our proposed method consistently addresses data redundancy across different datasets and architectures.
@@
Compared with MLLM-specialized selection methods, DataTailor exceeds SOTA for overall benchmarks. Specifically, DataTailor achieves an average of 100.1% relative performance, outperforming TIVE [43] (94.6%) and the latest ICONS [61] (98.8%). Notably, when increasing data selection ratio to 15%, DataTailor surpasses the performance of the full tuning of LLaVA-v1.5-7B (103.0% for MLLM benchmarks and 101.3% for total benchmarks). This indicates that a small amount of high-quality data is more crucial than a large volume of low-quality data for enhancing MLLMs, which truly exemplifies “Less is More”.
@@
mPLUG-264k [67]
   Full Data (100%) - 1243.4 34.3 26.9 41.1 100.0%
   Random (5%) - 1183.2 33.9 26.6 40.6 97.9%
   TIVE [43] (5%) LLaVA-vicuna-7B 1177.1 34.0 26.9 40.2 97.9%
 DataTailor (5%) LLaVA-vicuna-7B 1260.0 34.5 27.8 41.9 101.8%
mPLUG-OWL-7B 1217.2 34.3 28.0 41.3 100.6%
Bunny-695k [22]
   Full Data (100%) - 1778.1 70.7 38.7 70.9 100.0%
   Random (5%) - 1578.8 64.4 36.7 70.1 93.4%
   TIVE [43] (5%) LLaVA-vicuna-7B 1542.4 61.7 34.7 68.4 90.0%
 DataTailor (5%) LLaVA-vicuna-7B 1582.8 65.3 37.1 75.8 96.0%
Bunny-Phi-3B 1599.8 63.3 37.3 74.4 95.2%